OpenStudy (thatgirl3390):

Inverse function of F(x)=3^x+1

OpenStudy (nincompoop):

can you solve for x?

OpenStudy (anonymous):

To find the inverse: Replace f(x) with y Switch x's and y's, so put x where y is and x where y is. Solve for y Replace y with f^-1(x)

OpenStudy (nincompoop):

change f(x) into y then solve for x

OpenStudy (anonymous):

Nincompoop meant solve for Y

OpenStudy (nincompoop):

I mean SOLVE FOR X

OpenStudy (anonymous):

Whoops I read it wrong

OpenStudy (nincompoop):

after you solve for x, you swap x and y then rewrite "y" into f^-1

OpenStudy (thatgirl3390):

thx guys

OpenStudy (anonymous):

F(x) can be considered to be the same thing as Y so if you switch it for y you have\[y = 3^x + 1 \]

OpenStudy (nincompoop):

people swap in the beginning, I swap at the end

OpenStudy (anonymous):

1) subtract 1 2) take the log (base 3)

OpenStudy (nincompoop):

for x as an exponent you're going to need to learn about logarithm

OpenStudy (anonymous):

math is fun? must be tripping

OpenStudy (nincompoop):

hahaha at least it is not purple math ... when it is the brain that turns purple after doing lots of math

OpenStudy (thatgirl3390):

what about f^-1(10)

OpenStudy (anonymous):

translate as "solve for \(x\)"\[10 = 3^x + 1\]

OpenStudy (anonymous):

subtract \(1\) and get \[9=3^x\]then the answer should be more or less obvious right?

OpenStudy (thatgirl3390):

wow thx

OpenStudy (anonymous):

yw did you get \(x=2\) ?

OpenStudy (anonymous):

since \[f(2)=3^2+1=10\] that means \[f^{-1}(10)=2\]